# PMGDA: A Preference-based Multiple Gradient Descent Algorithm

### 14 Feb 2024 v1

@@ III-B The Predict-and-Correct Framework
Our proposed algorithm differs from Gebken et al. [27], which is primarily conceptual and not directly suited for large-scale neural networks. Their assumption of an available projection operator (${\bm{\theta}}\leftarrow\text{proj}({\bm{\theta}}-\eta{\bm{d}})$) is unrealistic in the context of neural networks. We have significantly modified this correction step in Problem (9) to enhance learning and investigated the prediction algorithm in its dual form. Consequently, our correction and prediction steps involve only $O(m)$ decision variables, compared to $O(n)$ in Gebken et al. [27], where $m$ is the number of objectives and $n$ is the number of decision variables. Given that $n\gg m$, our approach is the first practical method for finding preference-based Pareto solutions in neural networks.
III-C Efficient Prediction Algorithm
To solve problem (Equation 8) more efficiently, we explore its dual form. We define the Lagrangian function as $\mathcal{L}(\alpha,{\bm{v}},{\bm{\mu}},\nu)=\alpha+\frac{1}{2}\|{\bm{v}}\|^{2}+\sum_{i=1}^{m}{\bm{\mu}}_{i}(\nabla{{\bm{L}}_{i}}^{T}{\bm{v}}-\alpha)+\nu(\nabla h^{T}{\bm{v}})$, where ${\bm{\mu}}_{i}\geq 0$. Setting the derivative $\frac{\partial\mathcal{L}}{\partial{\bm{v}}}=0$ leads to $v=-\sum_{i=1}^{m}{\bm{\mu}}_{i}\nabla{\bm{L}}_{i}-\nu\nabla h$. Additionally, taking $\frac{\partial\mathcal{L}}{\partial\alpha}=0$ results in $\mu\in\Delta^{m-1}$. Consequently, the dual problem of Equation 8 is formulated as:
$\min_{({\bm{\mu}},\nu)\in\Delta^{m-1}\times\mathbb{R}}\left\|\sum_{i=1}^{m}{\bm{\mu}}_{i}\nabla{\bm{L}}_{i}+\nu\nabla h\right\|.$ (11)
The resulting direction ${\bm{d}}=-\sum_{i=1}^{m}{\bm{\mu}}_{i}^{*}\nabla{\bm{L}}_{i}-\nu^{*}\nabla h$, where ${\bm{\mu}}_{i}^{*}$ and $\nu^{*}$ are solutions to Problem Equation (11). It is noteworthy that Equation (11) involves only $m+1$ decision variables and constitutes a semi-positive definite programming problem, which can be solved very efficiently compared with it original version (Equation 8).
III-D Efficient Correction Algorithm
@@
$^{\bm{d}}=-\sum_{i=1}^{m}{\bm{\mu}}_{i}\widehat{\nabla{\bm{L}}_{i}}-{\bm{\mu}}_{m+1}\widehat{\nabla h}.$ (12)
Here, the notation $\hat{(\cdot)}$ is used to denote vector normalization, where $\hat{\mathbf{x}}=\frac{\mathbf{x}}{\|\mathbf{x}\|}$. Without loss of generality, we can assume that $(\mathbf{\mu}_{1},\ldots,\mathbf{\mu}_{m+1})\in\Delta^{m}$, as we can adjust the learning rate $\eta$ in Equation (10) to scale $\mathbf{d}$.
@@
$\left\|\sum_{i=1}^{m+1}{\bm{\mu}}_{i}{\bm{v}}_{i}\right\|\leq\sum_{i=1}^{m+1}\left\|{\bm{\mu}}_{i}{\bm{v}}_{i}\right\|\leq\sum_{i=1}^{m+1}|{\bm{\mu}}_{i}|=1,$
when ${\bm{v}}_{i}$ is a unit vector and $\mu\in\Delta_{m+1}$. We then show that when ${\bm{\theta}}_{k}$ is not Pareto stationary, there does not exist constants ${\bm{\mu}}_{1},\ldots,{\bm{\mu}}_{m+1}$ such that ${\bm{v}}=\sum_{i=1}^{m}{\bm{\mu}}_{i}\widehat{\nabla{\bm{L}}_{i}}+{\bm{\mu}}_{m+1}\widehat{\nabla h}=0$, i.e., $\left\|{\bm{v}}\right\|>0$. $\diamondsuit$
@@
$\nabla h^{T}\bm{v}\leq-\sigma\left\|\nabla h\right\|\cdot\left\|\bm{v}\right\|.$ (13)
@@
$\nabla h^{T}\bm{v}\leq-\sigma\left\|\nabla h\right\|.$ (14)
@@
Let $G$ (shape is $n\times(m+1)$) be the Jacobian matrix of the function $[{\bm{L}}({\bm{\theta}}),h({\bm{\theta}})]$. $G$ is formed by concatenating the columns $\nabla{\bm{L}}_{1}$, $\nabla{\bm{L}}_{2}$, …, $\nabla{\bm{L}}_{m}$, and $\nabla h$.
$G=[\nabla{\bm{L}}_{1},\ldots,\nabla{\bm{L}}_{m},\nabla h].$
Let $\hat{G}$ be the normalized augmented Jacobian matrix, where each column is normalized. Then, with Equation (14) and Lemma 1, the original optimization problem in the correction step (Equation (9)) can be simplified to the following easy-to-solve linear problem:
@@
$s.t.$ $\;(-G^{T}\hat{G}){\bm{\mu}}\leq\bm{b},$
where $\mathbf{b}$ is the vector for the linear constraint which is defined as $\mathbf{b}=[\underbrace{\alpha,...,\alpha}_{m+1},-\sigma\|\nabla h\|]^{T}$. We use $\tilde{\mathbf{\mu}}$ to denote the final decision variable, $\tilde{\mathbf{\mu}}=[\mathbf{\mu},\alpha]$. We can write Problem (15) in a compact matrix form that can be directly solved using the ‘cvxpy’ library [31]. The formulation is as follows:
@@ III-E Reuse of Gradients
For simple problems, obtaining the gradient of $\nabla h$ directly is cheap. However, for large-scale problems, such as neural networks, computing gradients becomes expensive. Note that $h({\bm{\theta}})$ is a function that only involves objectives ${\bm{L}}_{i}$’s. Once all the gradients $\nabla{\bm{L}}_{i}$’s are obtained, the gradient of the constraint function $\nabla h$ can be estimated using the chain rule:
$\nabla h=\sum_{i=1}^{m}\left(\frac{\partial h}{\partial{\bm{L}}_{i}}\cdot\nabla{\bm{L}}_{i}\right),$ (17)
where the first term $\frac{\partial h}{\partial{\bm{L}}_{i}}$ are just scalars.
For the specific case when $m=2$ (two-objective problem) and $h(\cdot)$ is defined by Equation (6), the partial derivatives $\frac{\partial h}{\partial{\bm{L}}_{i}}$ can be calculated using the following closed-form:
$\begin{cases}\frac{\partial h}{\partial{\bm{L}}_{1}}=\left(-\frac{{\bm{\lambda}}_{2}}{\|{\bm{\lambda}}\|},\frac{{\bm{\lambda}}_{1}}{\|{\bm{\lambda}}\|}\right),&\text{if }\frac{{\bm{L}}_{1}({\bm{\theta}})}{{\bm{\lambda}}_{1}}<\frac{{\bm{L}}_{2}({\bm{\theta}})}{{\bm{\lambda}}_{2}},\\ \frac{\partial h}{\partial{\bm{L}}_{2}}=\left(\frac{{\bm{\lambda}}_{2}}{\|{\bm{\lambda}}\|},-\frac{{\bm{\lambda}}_{1}}{\|{\bm{\lambda}}\|}\right),&\text{otherwise},\end{cases}$ (18)
@@
For other preference functions $h({\bm{\theta}})$ which may not have an analytical gradient expression, it is possible to perform a lightweight backward propagation to obtain the values of $\frac{\partial h}{\partial{\bm{L}}_{i}}$.
@@
We now demonstrate that the proposed PMGDA can be implemented even more efficiently through a dynamic weight adjustment mechanism. The final direction vector ${\bm{d}}$ can be succinctly expressed using the following equations, which is only a linear combination of gradients $\nabla{\bm{L}}_{i}$’s.
${\bm{d}}$ $=-\left(\sum_{i=1}^{m}{\bm{\mu}}^{*}_{i}\widehat{\nabla{\bm{L}}_{i}}+{\bm{\mu}}_{m+1}^{*}\widehat{\nabla h}\right)$ (19)
(Substitute the chain rule (Equation (17)))
$=-\left(\sum_{i=1}^{m}\frac{{\bm{\mu}}^{*}_{i}}{\left\|\nabla{\bm{L}}_{i}\right\|}\nabla{\bm{L}}_{i}+\frac{{\bm{\mu}}_{m+1}^{*}}{\left\|\nabla h\right\|}\sum_{i=1}^{m}\frac{\partial h}{\partial{\bm{L}}_{i}}\nabla{\bm{L}}_{i}\right)$
$=-\left(\sum_{i=1}^{m}\left(\frac{{\bm{\mu}}^{*}_{i}}{\left\|\nabla{\bm{L}}_{i}\right\|}+\frac{{\bm{\mu}}_{m+1}^{*}}{\left\|\nabla h\right\|}\frac{\partial h}{\partial{\bm{L}}_{i}}\right)\nabla{\bm{L}}_{i}\right).$
$=-\left(\sum_{i=1}^{m}{\bm{\mu}}_{i}^{\prime}\nabla{\bm{L}}_{i}\right)\quad{\bm{\mu}}_{i}^{\prime}:=\left(\frac{{\bm{\mu}}^{*}_{i}}{\left\|\nabla{\bm{L}}_{i}\right\|}+\frac{{\bm{\mu}}_{m+1}^{*}}{\left\|\nabla h\right\|}\frac{\partial h}{\partial{\bm{L}}_{i}}\right).$
III-F Algorithm and Complexity Analysis
@@
             Find $\mu^{*}$ by solving the linear programming problem Equation (16).
             Calculate the dynamic weight $\mu^{\prime}$ by Equation (19).
             ${\bm{d}}=-\sum_{i=1}^{m}\mu^{\prime}_{i}\nabla{\bm{L}}_{i}$.
@@
Algorithm 1 The PMGD Algorithm
We summarize our practical method as Algorithm 1 and call it Preference Multiple Gradient Descent Algorithm (PMGDA). The stopping condition used in our experiments is that $\|d\|\leq\epsilon$ and $h({\bm{\theta}}_{k})\leq\delta$. This condition is met when $d=0$ and $h({\bm{\theta}}_{k})=0$, indicating that the solution meets user’s demand.
The main time consumption of PMGDA in the ‘correction’ step is the calculation of $G^{T}\hat{G}$, which has a complexity of $\mathcal{O}(({m+1})^{2}n)$. This calculation time dominates the optimization time, which has a ${\mathcal{O}}^{*}({(2m+3)}^{2.37})$ complexity. According to out empirical results, this step is converged in only several iterarions. In the ‘prediction’ step, the running time of Pure MGDA [19, 20] is $\mathcal{O}(m^{2}n)$. It is worth noting that MGDA does not guarantee the position of the final Pareto solution. On the other hand, our method can generate a specific solution satisfying a user’s demand without significantly increasing complexity, which is advantageous. The complexity of the number of variables is linear, while the complexity with respect to the objective is quadratic.
@@ III-G Multi-objective Reinforcement Learning (MORL) Task
In this section, we demonstrate the adaptability of our proposed framework to MORL tasks, which are very practical real-world problems. In MORL, each state-action pair $(s,a)$ is associated with $m$ conflicting rewards, denoted as ${\bm{r}}_{1}(s,a),\ldots,{\bm{r}}_{m}(s,a)$. The objective in MORL is to optimize $m$ cumulative returns, each defined by:
${\bm{L}}_{i}({\bm{\theta}})=\mathbb{E}_{\pi\sim\pi_{{\bm{\theta}}}({\bm{\lambda}})}\left[\sum_{t=1}^{T}\gamma^{t}{\bm{r}}_{i}(s_{t},a_{t})\right],\forall i\in[m],$ (20)
where $\pi_{{\bm{\theta}}}({\bm{\lambda}})$ represents a policy network parameterized by ${\bm{\theta}}$, mapping states to actions. Here, $10<\gamma<$ 1 is the discount factor, and $T$ signifies the total time steps. The gradient of ${\bm{L}}_{i}({\bm{\theta}})$ can be estimated through interaction with the environment. We illustrate this using the most direct policy gradient method [33], shown as:
$\nabla{\bm{L}}_{i}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{{\bm{\theta}}}\log\pi_{{\bm{\theta}}}(a_{t}|s_{t})(\bm{G}_{i}(s_{t})-\bm{b}_{i}(s_{t}))\right],$ (21)
where $\bm{G}_{i}(s_{t})=\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}{\bm{r}}_{t^{\prime}}$ represents the discounted cumulative reward from timestep $t$ to $T$, and $\bm{b}_{i}(s_{t})$ is a baseline function. Practically, $\bm{b}_{i}(s_{t})$ is approximated by a value network $\bm{V}_{\bm{\phi}_{i}}(s)$, and $\nabla{\bm{L}}_{i}$ is estimated using $N$ trajectory samples:
$\nabla{\bm{L}}_{i}\approx\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{T-1}\nabla_{{\bm{\theta}}}\log\pi_{{\bm{\theta}}}(a_{t}|s_{t})(\bm{G}_{i}(s_{t})-\bm{V}_{\bm{\phi}_{i}}(s_{t})).$ (22)
@@
Secondly, MORL involves maximization rather than minimization, hence, the constraint formulation needs to be adjusted accordingly. The constraint $\nabla{\bm{L}}_{i}^{T}\bm{v}\leq\alpha,\;i\in[m]$ should be altered to a ‘greater than’ condition in Equations (8) and (9).

### 16 Feb 2024 v2

@@ III-B The Predict-and-Correct Framework
Our proposed algorithm differs from Gebken et al. [27], which is primarily conceptual and not directly suited for large-scale neural networks. Their assumption of an available projection operator (${\bm{\theta}}\leftarrow\text{proj}({\bm{\theta}}-\eta{\bm{d}})$) is unrealistic in the context of neural networks. We have significantly modified this correction step in Problem (9) to enhance learning and investigated the prediction algorithm in its dual form. Consequently, our correction and prediction steps involve only $O(m)$ decision variables, compared to $O(n)$ in Gebken et al. [27]’s original work, where $m$ is the number of objectives and $n$ is the number of decision variables. Given that $n\gg m$, our approach is the first practical method for finding preference-based Pareto solutions in neural networks.
III-C An Efficient Prediction Algorithm
To solve problem (Equation 8) more efficiently, we explore its dual form. We define the Lagrangian function as $\mathcal{L}(\alpha,{\bm{v}},{\bm{\mu}},\nu)=\alpha+\frac{1}{2}\|{\bm{v}}\|^{2}+\sum_{i=1}^{m}\mu_{i}(\nabla{L_{i}}^{\top}{\bm{v}}-\alpha)+\nu(\nabla h^{\top}{\bm{v}})$, where $\mu_{i}\geq 0$. Setting the derivative $\frac{\partial\mathcal{L}}{\partial{\bm{v}}}=\bm{0}$ leads to ${\bm{v}}=-\sum_{i=1}^{m}\mu_{i}\nabla L_{i}-\nu\nabla h$. Additionally, taking $\frac{\partial\mathcal{L}}{\partial\alpha}=0$ results in ${\bm{\mu}}\in\Delta^{m-1}$. Consequently, the dual problem of Equation 8 is formulated as:
$\min_{({\bm{\mu}},\nu)\in\Delta^{m-1}\times\mathbb{R}}\left\|\sum_{i=1}^{m}\mu_{i}\nabla L_{i}+\nu\nabla h\right\|.$ (11)
The resulting optimal direction ${\bm{d}}=-\sum_{i=1}^{m}\mu_{i}^{*}\nabla L_{i}-\nu^{*}\nabla h$, where $\mu_{i}^{*}$ and $\nu^{*}$ are solutions to problem of Equation (11). It is noteworthy that Equation 11 only involves $m+1$ decision variables and constitutes a semi-positive definite programming problem, which can be solved very efficiently compared with it original problem (Equation 8).
III-D An Efficient Correction Algorithm
@@
$^{\bm{d}}=-\sum_{i=1}^{m}\mu_{i}\widehat{\nabla L_{i}}-\mu_{m+1}\widehat{\nabla h}.$ (12)
Here, the notation $\hat{(\cdot)}$ is used to denote vector normalization, where $\hat{{\bm{x}}}=\frac{{\bm{x}}}{\|{\bm{x}}\|}$. Without loss of generality, we can assume that $(\mu_{1},\ldots,\mu_{m+1})\in\Delta^{m}$, as we can adjust the learning rate $\eta$ in Equation (10) to scale the norm of updating direction ${\bm{d}}$.
@@
$\left\|\sum_{i=1}^{m+1}\mu_{i}v_{i}\right\|\leq\sum_{i=1}^{m+1}\left\|\mu_{i}v_{i}\right\|\leq\sum_{i=1}^{m+1}|\mu_{i}|=1,$
when $v_{i}$ is a unit vector and $\mu\in\Delta_{m+1}$. We then show that when ${\bm{\theta}}_{k}$ is not Pareto stationary, there does not exist constants $\mu_{1},\ldots,\mu_{m+1}$ such that ${\bm{v}}=\sum_{i=1}^{m}\mu_{i}\widehat{\nabla L_{i}}+\mu_{m+1}\widehat{\nabla h}=\bm{0}$, i.e., $\left\|{\bm{v}}\right\|>0$. $\diamondsuit$
@@
$\nabla h^{\top}{\bm{v}}\leq-\sigma\left\|\nabla h\right\|\cdot\left\|{\bm{v}}\right\|.$ (13)
@@
$\nabla h^{\top}{\bm{v}}\leq-\sigma\left\|\nabla h\right\|.$ (14)
@@
Let ${\bm{G}}$ (shape is $n\times(m+1)$) be the Jacobian matrix of the function $[{\bm{L}}({\bm{\theta}}),h({\bm{\theta}})]$. ${\bm{G}}$ is formed by concatenating the columns $\nabla L_{1}$, $\nabla L_{2}$, …, $\nabla L_{m}$, and $\nabla h$.
${\bm{G}}=[\nabla L_{1},\ldots,\nabla L_{m},\nabla h].$
Let $\hat{{\bm{G}}}$ be the normalized augmented Jacobian matrix, where each column is normalized. Then, with Equation (14) and Lemma 1, the original optimization problem in the correction step (Equation (9)) can be simplified to the following easy-to-solve linear problem:
@@
$s.t.$ $\;(-{\bm{G}}^{\top}\hat{{\bm{G}}}){\bm{\mu}}\leq{\bm{b}},$
where ${\bm{b}}$ is the vector for the linear constraint which is defined as ${\bm{b}}=[\underbrace{\alpha,...,\alpha}_{m+1},-\sigma\|\nabla h\|]^{\top}$. We use $\tilde{{\bm{\mu}}}$ to denote the final decision variable, $\tilde{{\bm{\mu}}}=[{\bm{\mu}},\alpha]$. In this way, Problem (15) can be written in a compact matrix form that can be directly solved using the ‘cvxpy’ library [31]. The formulation is as follows:
@@ III-E Reuse of Gradients
For simple problems, obtaining the gradient of $\nabla h$ directly is cheap. However, for large-scale problems, such as neural networks, computing gradients becomes expensive. Note that $h({\bm{\theta}})$ is a function that only involves objectives $L_{i}$’s. Once all the gradients $\nabla L_{i}$’s are obtained, the gradient of the constraint function $\nabla h$ can be estimated using the chain rule:
$\nabla h=\sum_{i=1}^{m}\left(\frac{\partial h}{\partial L_{i}}\cdot\nabla L_{i}\right),$ (17)
where the first term $\frac{\partial h}{\partial L_{i}}$ are just scalars.
For the specific case when $m=2$ (two-objective problem) and $h(\cdot)$ is defined by Equation (6), the partial derivatives $\frac{\partial h}{\partial L_{i}}$ can be calculated using the following closed-form:
$\begin{cases}\frac{\partial h}{\partial L_{1}}=\left(-\frac{\lambda_{2}}{\|{\bm{\lambda}}\|},\frac{\lambda_{1}}{\|{\bm{\lambda}}\|}\right),&\text{if }\frac{L_{1}({\bm{\theta}})}{\lambda_{1}}<\frac{L_{2}({\bm{\theta}})}{\lambda_{2}},\\ \frac{\partial h}{\partial L_{2}}=\left(\frac{\lambda_{2}}{\|{\bm{\lambda}}\|},-\frac{\lambda_{1}}{\|{\bm{\lambda}}\|}\right),&\text{otherwise},\end{cases}$ (18)
@@
For other preference functions $h({\bm{\theta}})$ which may not have an analytical gradient expression, it is possible to perform a lightweight backward propagation to obtain the values of $\frac{\partial h}{\partial L_{i}}$.
@@
We now demonstrate that the proposed PMGDA can be implemented even more efficiently through a dynamic weight adjustment mechanism. The final direction vector ${\bm{d}}$ can be succinctly expressed using the following equations, which is only a linear combination of gradients $\nabla L_{i}$’s.
${\bm{d}}$ $=-\left(\sum_{i=1}^{m}\mu_{i}^{*}\widehat{\nabla L_{i}}+\mu_{m+1}^{*}\widehat{\nabla h}\right)$ (19)
(Substitute the chain rule (Equation (17)) to the above equation.)
$=-\left(\sum_{i=1}^{m}\frac{\mu_{i}^{*}}{\left\|\nabla L_{i}\right\|}\nabla L_{i}+\frac{\mu_{m+1}^{*}}{\left\|\nabla h\right\|}\sum_{i=1}^{m}\frac{\partial h}{\partial L_{i}}\nabla L_{i}\right)$
$=-\left(\sum_{i=1}^{m}\left(\frac{{\bm{\mu}}^{*}_{i}}{\left\|\nabla L_{i}\right\|}+\frac{\mu_{m+1}^{*}}{\left\|\nabla h\right\|}\frac{\partial h}{\partial L_{i}}\right)\nabla L_{i}\right).$
$=-\left(\sum_{i=1}^{m}\mu_{i}^{\prime}\nabla L_{i}\right)\quad\mu_{i}^{\prime}:=\left(\frac{\mu_{i}^{*}}{\left\|\nabla L_{i}\right\|}+\frac{\mu_{m+1}^{*}}{\left\|\nabla h\right\|}\frac{\partial h}{\partial L_{i}}\right).$
III-F Algorithm and its Complexity Analysis
@@
             Solve the coefficient ${\bm{\mu}}^{*}$ in linear programming problem (Equation (16)).
             Calculate the dynamic weight ${\bm{\mu}}^{\prime}$ by Equation (19).
             ${\bm{d}}=-\sum_{i=1}^{m}\mu_{i}^{\prime}\nabla L_{i}$.
@@
Algorithm 1 The Preference based Multiple Gradient Descent Algorithm (PMGDA)
We summarize our practical method as Algorithm 1 and call it PMGDA. The stopping condition used in our experiments is that $\|d\|\leq\epsilon$ and $h({\bm{\theta}}_{k})\leq\delta$. This condition is met when $d=0$ and $h({\bm{\theta}}_{k})=0$, indicating that the solution meets decision maker’s demand.
The main time consumption of PMGDA in the ‘correction’ step is the calculation of ${\bm{G}}^{\top}\hat{{\bm{G}}}$, which has a complexity of $\mathcal{O}(({m+1})^{2}n)$. This calculation time dominates the optimization time, which has a ${\mathcal{O}}^{*}({(2m+3)}^{2.37})$ complexity. According to out empirical results, this step is converged in only several iterations. In the ‘prediction’ step, the running time of Pure MGDA [19, 20] is $\mathcal{O}(m^{2}n)$. It is worth noting that MGDA does not guarantee the position its searched Pareto solution. On the other hand, it is advantageous our method can generate a specific solution satisfying a decision maker’s demand without significantly increasing complexity. The complexity of the number of variables is linear, while the complexity with respect to the objective is quadratic.
@@ III-G Multi-objective Reinforcement Learning (MORL) Task
In this section, we demonstrate the adaptability of our proposed framework to MORL tasks, which are very practical real-world problems. In MORL, each state-action pair $(s,a)$ is associated with $m$ conflicting rewards, denoted as $r_{1}(s,a),\ldots,r_{m}(s,a)$. The objective in MORL is to optimize $m$ cumulative returns, each defined by:
$L_{i}({\bm{\theta}})=\mathbb{E}_{\pi\sim\pi_{{\bm{\theta}}}({\bm{\lambda}})}\left[\sum_{t=1}^{T}\gamma^{\top}r_{i}(s_{t},a_{t})\right],\forall i\in[m],$ (20)
where $\pi_{{\bm{\theta}}}({\bm{\lambda}})$ represents a policy network parameterized by ${\bm{\theta}}$, mapping states to actions. Here, $10<\gamma<$ 1 is the discount factor, and $T$ signifies the total time steps. The gradient of $L_{i}({\bm{\theta}})$ can be estimated through interaction with the environment. We illustrate this using the most direct policy gradient method [33], shown as:
$\nabla L_{i}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{{\bm{\theta}}}\log\pi_{{\bm{\theta}}}(a_{t}|s_{t})(\bm{G}_{i}(s_{t})-{\bm{b}}_{i}(s_{t}))\right],$ (21)
where $\bm{G}_{i}(s_{t})=\sum_{t^{\prime}=t}^{\top}\gamma^{t^{\prime}-t}r_{t^{\prime},i}$ represents the discounted cumulative reward from timestep $t$ to $T$, and ${\bm{b}}_{i}(s_{t})$ is a baseline function. Practically, ${\bm{b}}_{i}(s_{t})$ is approximated by a value network ${\bm{v}}_{\bm{\phi}_{i}}(s)$, and $\nabla L_{i}$ is estimated using $N$ trajectory samples:
$\nabla L_{i}\approx\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{T-1}\nabla_{{\bm{\theta}}}\log\pi_{{\bm{\theta}}}(a_{t}|s_{t})(\bm{G}_{i}(s_{t})-{\bm{v}}_{\bm{\phi}_{i}}(s_{t})).$ (22)
@@
Secondly, MORL involves maximization rather than minimization, hence, the constraint formulation needs to be adjusted accordingly. The constraint $\nabla L_{i}^{\top}{\bm{v}}\leq\alpha,\;i\in[m]$ should be altered to a ‘greater than’ condition in Equations (8) and (9).